# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v2

@@ Appendix C Normalization
$\boldsymbol{a}^{\prime(S2)}_{j}[t]=\begin{cases}\boldsymbol{a}^{\prime(S2)}_{j}[t],&\text{ if }\boldsymbol{a}^{\prime(S2)}_{j}[t]\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\prime(S2)}_{j}[t]-\lambda),&\text{otherwise}.\end{cases}$ (9)
@@
According to this formula, $\boldsymbol{a}^{\prime}$ grows undiminished with $\boldsymbol{a}^{\prime}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$. (In our experiments we found the value $\lambda=1.3\cdot\frac{h^{(W2)}+w^{(W2)}}{2}$ to work well). This saturation helps to mitigate imbalances between net fragments with differing numbers of participating neurons. In the next step, we introduce the normalization
@@
$\boldsymbol{a}^{\text{norm }(S2)}_{c,i}[t]=\frac{\boldsymbol{a}^{\prime(S2)}_{c,i}[t]}{\max_{i^{\prime}}\boldsymbol{a}^{\prime(S2)}_{c,i^{\prime}}[t]}$ (10)
@@
This normalization confines activations to the interval $(0,1)$, where neurons receiving maximal support are mapped to $1$, while those with lesser support are mapped to correspondingly smaller values. This normalization relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later, when recurrent links of often co-active neurons have grown, to suppress those neurons that are part of accidental or noise patterns.
Appendix D Alternative Neuron Selection
Only one of a set of alternative neurons is allowed to be active. The selection process is determined by evaluating the correlation between feature patches and alternative kernel filters. Initially, patches matching the size of the feature kernel are extracted from every position in the input $(\boldsymbol{y}^{(S1)};\boldsymbol{y}^{(S2)}[t-1])$. Subsequently, the correlation of these patches with the feature kernels is computed, and at each location, the alternative kernel exhibiting the highest correlation is chosen. Consequently, multiple features may be active at each location, but only one of the competing feature channels — the one best matching a given input patch — achieves activation. At initialization (see Appendix A), where all alternative channels fit equally well to a given local pattern, the first channel is selected. However, after applying a Hebbian update, this channel starts specializing in recognizing a specific local pattern. Consequently, the activation of subsequent local patterns may either trigger the same channel if they have a high correlation or activate another channel better suited to recognizing the new pattern.
Appendix E Dataset
The dataset comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles).
@@ Appendix E Dataset
The images are generated when required and may vary from one epoch to another. During each training cycle, we randomly generate $300$ image instances (each line about $5$ times). During evaluation, we sample each of the $59$ distinct lines exactly once. In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
@@
To demonstrate the generalization capability of CNA areas, we generate validation samples that differ from the ones used during training. Specifically, we produce binary images measuring ($32\times 32$) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
@@
Appendix F Feature Extraction
The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location. We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 10, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal). This approach is motivated by the observation that, given our specific elementary feature types (cf. Figure 10), straight lines, when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons). The use of such manually designed filters allows a more straightforward interpretation of the results. While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
These filters move across the image with a step size (‘stride’) of $1$ (assuming 00 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size. Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$. The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels. We convert the floating-point output of the convolutional weights to binary activations using binary neurons $\mathrm{B}()$ with a bias of $b^{(S1)}=0.5$, i.e. neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq. 2).
Appendix G Parameters
We train the model with a learning rate of $\alpha=0.2$ for $100$ training cycles (epochs). The model uses $n_{a}=10$ alternative neurons with recurrent connections spanning a kernel of $11\times 1$ 11 neurons. Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix F), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot n_{a}=40$. The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
Appendix H Autoencoder Parameter
@@
The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model. It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ layers. The decoder mirrors this structure, using transposed convolutional layers with $256$, $128$, $64$, and $32$ channels, respectively. ReLU activations are applied after each layer. Consequently, the autoencoder is a relatively large model with $777$k parameters, utilizing up to $256$ channels, whereas the CNA model operates with only $44$ channels.
@@
Appendix I Quantifying Noise Filtration
@@
Figure 11: System performance as function of the noise level. First row: noise reduction rate; second row: recall; third row: precision. Columns differ in the inhibition coefficient $\gamma$ that models inhibition. The colors within each plot represent different activation biases $b^{(S2)}$. The dotted red line is the autoencoder, serving as baseline.
Figure 11 displays noise reduction rates, recall, and precision across different parameter configurations for different levels of added noise. The metrics demonstrate that increasing the inhibition coefficient or the activation bias results in enhanced noise filtration.
@@
Figure 12: Reconstruction of partial occlusion. First row: feature reconstruction rate; second row: recall; third row: precision. Columns are distinguished by the inhibition coefficient $\gamma$, and line colors within plots distinguish activation bias $b^{(S2)}$. The dotted red line is the autoencoder model.
Figure 12 displays the feature reconstruction rate, precision, and recall for removed line fragments of different lengths. Having a lower activation bias or inhibition coefficient $\gamma$ leads to a better reconstruction of removed patterns, as these settings allow neurons with less support to activate.

### v3

@@ Appendix B Normalization
$\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}=\begin{cases}\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t},&\text{ if }\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}\leq\lambda\\ \lambda-\frac{1}{2}(\boldsymbol{a}^{\mathcal{M}(S2)}_{c,t}-\lambda),&\text{otherwise}.\end{cases}$ (21)
@@
According to this formula, $\boldsymbol{a}^{\mathcal{M}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{M}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope $-1/2$. (In our experiments we found the value $\lambda=1.3\cdot\frac{h+w}{2}$ to work well). This saturation helps to mitigate imbalances between net fragments with differing numbers of participating neurons. In the next step, we introduce the normalization
@@
$\boldsymbol{a}^{\text{norm }(S2)}_{c,i,t}=\max\left(\frac{\boldsymbol{a}^{\mathcal{M}(S2)}_{c,i,t}}{\max_{i^{\prime}}\boldsymbol{a}^{\mathcal{M}(S2)}_{c,i^{\prime},t}},0\right)$ (22)
@@
Here, $i$ refers to spatial locations within the image, i.e., $i=(h,w)$ where $h\in H,w\in W$. This normalization confines activations to the interval $[0,1]$, where neurons receiving maximal support are mapped to $1$, while those with less support are mapped to correspondingly smaller values. This normalization is relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later when recurrent weights between correlated patterns have been learned, to suppress those neurons that are part of noise patterns.
Appendix C Single Weight-Kernel Implementation
So far, we introduced the weights $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ in $S2$ as two independent convolutional kernels. For simplicity in our algorithm, and since we use the same kernel size for the forward and recurrent connections, we stack the two kernels as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}^{(S1)}$ and $\boldsymbol{y}^{(S2)}_{t-1}$, leading to the equivalent formulation of eq. 16:
$\boldsymbol{a}^{\mathcal{M}(S2)}_{c_{\text{out}},t}=\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(S2)}_{c_{\text{out}},c_{\text{in}}}\star\left(\boldsymbol{y}_{c_{\text{in}}}^{(S1)};\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)$ (23)
Appendix D Dataset
@@
The dataset used for training comprises binary images measuring ($32\times 32$) pixels, each sample depicting a straight line going through the image center, starting and ending $2$ pixels from the image boundary (leading to $59$ distinct lines of different angles). The images are generated when required and may vary from one epoch to another. During each training cycle, we randomly generate $300$ image instances (each line about $5$ times). During evaluation, we sample each of the $59$ distinct lines exactly once. In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
@@
To demonstrate the generalization capability of CNA, we generate validation samples that differ from the ones used during training. Specifically, we produce binary images measuring ($32\times 32$) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
@@
Appendix E Parameters
We train the model with a learning rate of $\eta=0.2$ for $100$ training cycles (epochs). The model uses $\kappa=10$ alternative neurons with recurrent connections spanning a kernel of $11\times 1$ 11 neurons. Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix A), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot\kappa=40$. The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
Appendix F Autoencoder Parameter
@@
The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model. It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ channels. The decoder mirrors this structure, using transposed convolutional layers with $256$, $128$, $64$, and $32$ channels, respectively. ReLU activations are applied after each layer. Consequently, the autoencoder is a relatively large model with $777$k parameters, utilizing up to $256$ channels, whereas the CNA model operates with only $44$ channels.
@@
Appendix G Quantifying Noise Filtration
@@
Figure 10: System performance as function of the noise level. First row: noise reduction rate; second row: recall; third row: precision. Columns differ in the attenuation coefficient $\gamma$. The colors within each plot represent different activation biases $b^{(S2)}$. The dotted red line is the autoencoder, serving as baseline.
Figure 10 displays noise reduction rates, recall, and precision across different parameter configurations for different levels of added noise. The metrics demonstrate that increasing the attenuation coefficient or the activation bias results in enhanced noise filtration.
@@
Figure 11: Reconstruction of partial occlusion. First row: feature reconstruction rate; second row: recall; third row: precision. Columns are distinguished by the attenuation coefficient $\gamma$, and line colors within plots distinguish activation bias $b^{(S2)}$. The dotted red line is the autoencoder model.
Figure 11 displays the feature reconstruction rate, precision, and recall for removed line fragments of different lengths. Having a lower activation bias or attenuation coefficient $\gamma$ leads to a better reconstruction of removed patterns, as these settings allow neurons with less support to activate.